# Raising the Bar of AI-generated Image Detection with CLIP

### v1

@@ Abstract
Aim of this work is to explore the potential of pre-trained vision-language models (VLMs) for universal detection of AI-generated images. We develop a lightweight detection strategy based on CLIP features and study its performance in a wide variety of challenging scenarios. We find that, unlike previous belief, it is neither necessary nor convenient to use a large domain-specific dataset for training. On the contrary, by using only a handful of example images from a single generative model, a CLIP-based detector exhibits a surprising generalization ability and high robustness across several different architectures, including recent commercial tools such as Dalle-3, Midjourney v5, and Firefly. We match the SoTA on in-distribution data, and improve largely above it in terms of generalization to out-of-distribution data (+6% in terms of AUC) and robustness to impaired/laundered data (+13%). Our project is available at https://grip-unina.github.io/ClipBased-SyntheticImageDetection/
@@ 1 Introduction
Synthetic images have by now left the research laboratories and are flooding the real world. The latest versions of popular image editing tools, such as Adobe Photoshop and Microsoft Paint, come with easy-to-use AI-powered generative tools that allow even novice users to edit and generate visual content at will. Thanks to diffusion-based generative models, it is not only the quality of the images generated that has improved enormously, but also the flexibility in using these tools. By issuing a few textual commands, it is possible to easily obtain the desired image. While this represents a great opportunity for visual art applications, it is also the paradise of disinformation professionals who can design their attacks with unprecedented power and flexibility [8, 5, 25]. And, of course, a nightmare for all those trying to combat the spread of fake news and orchestrated disinformation campaigns: journalists, fact-checkers, law enforcement agencies, governments. Therefore, there is high demand for automatic tools that help establishing the authenticity of a media asset.
@@
Figure 1: Area Under ROC Curve (AUC %) on unseen synthetic generators ($x$-axis) and on post-processed data ($y$-axis). The first number measures the generalization ability of the detector, the second its robustness to possible impairments. Circle area proportional to training set size. The performance is measured over 18 different synthetic models. Our CLIP-based detector largely outperforms all SOTA methods with very limited training data.
@@
It is well known that each synthesis model leaves its own peculiar traces in all generated images, subtle traces that give rise to so called artificial fingerprints, that can be exploited for forensic analyses [49, 77]. This was first observed for methods based on generative adversarial networks (GAN) [79, 22] but holds with obvious differences for all generation approaches, including the more recent diffusion models (DM) [13]. One of the main challenges for current forensic detectors is the ability to generalize to new and unseen generative methods. Indeed, the in-distribution (ID) scenario, with perfectly aligned training and test set, is rarely met in practice. Test images are often generated by new approaches, unseen architectures, or even known architectures re-trained under different conditions, all different flavors of the out-of-distribution (OOD) scenario. In addition, most of the times, images are downloaded from some social networks, where they undergo transformations such as compression and resizing (even multiple times), processes that tend to wash out the tiny traces so precious for detection.
In [71], it is shown that training diversity and intense augmentation are crucial for generalization. A ResNet-50 detector is trained on a single but very diverse dataset of about 360k ProGAN images [40]. Results on OOD images generated by other GAN-based generators turned out to be surprisingly good. The authors themselves, however, suggest that structural similarities between GAN-based generators are the likely reason for such good performance. Indeed, results are not equally good on images generated by recent diffusion-based methods, which present somewhat different generation artifacts [14].
@@
Generalizing to OOD data is a major issue in deep learning and has been the object of intense research in latest years. In this context, the advent of large pre-trained vision-language models has brought about a number of new solutions and exciting results. These models have been shown to be excellent zero-shot and few-shot learners in many diverse applications, such as image classification [59, 78], detection [26, 53] and segmentation [74, 81]. Recently, there have been attempts to exploit the power of VLMs to detect synthetic images [66, 2, 56]. For example, in [56] CLIP (Contrastive Language-Image Pre-Training) [59] is used as a pre-trained feature extractor, and a classifier is then trained on the 360k-image ProGAN dataset proposed in [71], with the same augmentation strategy. With respect to [71] based on Resnet50, the performance improves significantly, especially on images generated by diffusion models, showing a very good generalization ability.
@@
In this work we explore in depth the potential of CLIP for image forensics, propose new decision strategies, and carry out an extensive experimental analysis in challenging real-world scenarios involving a large number of generative models. We find that CLIP-based methods show impressive generalization ability and very good robustness, resulting in large performance improvements compared to the state of the art. To obtain such results we avoid any intensive training on domain-specific data, which may introduce unwanted biases and undermine the descriptive power of CLIP features. Instead, we rely on a small number of paired real/fake images with the same textual description and use their CLIP features to model the decision space. Synthetic images come from a single generator but work well with all other tested data sources. The top performance is obtained with just 1000 to 10000 paired images. Moreover, only minor decays are observed when this number reduces to 100 or even 10 (see Fig. 1).
In summary, the key contributions of this work are as follows:
@@
  • We carry out a large set of experiments on diverse synthetic generators and in very challenging conditions and are able to achieve the best performance (on average). Our experiments make clear that the features extracted are at least partially orthogonal to those explored by previous methods mostly oriented to extract low-level features.
@@ 2 Related work
There is an extensive literature on synthetic image detection, mostly concerning images created by generative adversarial networks (GANs) and, more recently, by diffusion models (DMs). Some methods look for visible errors, such as asymmetries in faces, wrong perspective, odd shadows [51, 28, 29]. However, with the rapid progress in image synthesis technology, this kind of problems are solved quickly and appear more and more rarely in modern generation methods. Hence, we will focus on methods that exploit “invisible” forensic traces, working either in the spatial or the frequency domain. Then we will briefly review the most recent approaches that exploit multimodal features.
@@
Spatial domain methods. Despite their high visual quality, synthetic images carry with them distinctive traces of the generation process that enable detection and even model identification [49, 77]. In fact, each generation model inserts a sort of digital fingerprint into all the images created, which depends on its architectural and training details. This fingerprint can be easily estimated, Given an adequate number of images generated by the target model, typically a few hundred, it is sufficient to average their noise residuals, extracted through simple denoisers [49] or more sophisticated deep learning-based methods [67, 47]. Detectors based on digital fingerprints can be regarded as “few-shot”, as they can deal with new models based on just a few example images. Other few-shot solutions have been proposed lately in this field [15, 50, 21, 36]. However, they all require some prior knowledge of the target models, even if limited to a few images, so they fit a strictly in-distribution (ID) scenario. In contrast, our work aims to generalize to new and previously unseen models, and therefore fits into an out-of-distribution (OOD) scenario.
@@
Good generalization is a key requirement of synthetic image detectors. Towards this end, it is important to increase diversity in training, which can be pursued by suitable augmentation, by including a large number of different categories [71], or even by model ensembling [48]. Other studies suggest to work on local patches [9] or combine global spatial information with local features [38]. In [33] it is shown that to preserve the subtle high-frequency forensic traces, one should avoid any downsampling in the first layers of the network. With the same aim, [69] proposes to work on noise residuals rather than the original images, by extracting the gradients with a pre-trained CNN. In [12] a systematic study is carried out on transferable forensic features that allows easier generalization.
@@
The above investigations and proposals, however, only consider GAN-based generators. Some very recent works extend the analyses to latest generation approaches. [14] shows that detectors designed for GAN images have a hard time generalizing to DM images, especially in the presence of common post-processing steps such as compression and resizing. Furthermore, they are unable to select an adequate decision threshold without the help of some calibration data from the model under test. On the other hand, [24] shows that it is possible to achieve decent performance by continuously re-training the detector on images from new generators, as long as the latter share some architectural components with the known ones. In [72], inspired to previous work on GAN images [1], a detector specifically tailored to DM images is proposed. Images are projected in a latent space and reconstructed with known models to study their features. While promising, the approach refers to specific architectures, hence an ID scenario, and has been tested on a limited set of diffusion generative models. In the present work, instead, we do not make any assumptions on the test data and include a large variety of both GAN and DM generators including OOD images, and realistic high-resolution images generated by recent diffusion models.
@@
Frequency domain methods. GAN-image artifacts are more easily spotted in the frequency domain and are clearly visible in the artificial fingerprint spectra [22, 31, 23]. In fact, they are caused by the up-sampling operations used in the generator, which give rise to regular spatial patterns and strong peaks in the Fourier domain. Interestingly, the awareness of such weakness has prompted the design of new architectures, like StyleGAN3 [43], which explicitly avoid aliasing and reduce the above mentioned peaks. In addition to such obvious artifacts, the spectral content of GAN images and real images is also known to differ significantly in the medium and high frequencies. Likewise, clear differences in the radial and angular spectral distributions of the DM and real images were observed [13, 75]. Furthermore, it is worth observing that DM images also may present spectral peaks (see Fig. 5).
Of course, frequency domain traces can be used to train simple detectors for ID images. Even more interestingly, some authors [79, 37] replicate and modify the synthesis process of known generators by introducing a series of small architectural changes in order to learn to manage also OOD test images. A major problem with spectral traces is that they can be easily weakened by laundering operations, both innocent (e.g., image resizing) and malicious counterforensic methods [20, 11]. Likewise, attackers can easily embed artificial spectral traces in a real image to make it appear as synthetic. This is a more general problem of low-level traces, which can be attacked in both ways and through a number of ingenious methods. We will show that our approach is indeed immune to such attacks and can handle a variety of distribution shift.
@@
Exploiting multimodal features. The introduction of large language models has elicited intense work in the vision community on trying to exploit large models trained using both images and text [80, 78]. In image forensics, however, traditionally oriented to the analysis of low-level features, there has been very little work in this direction. Only a few papers [66, 2, 56] have tried exploiting pre-trained visual models for the detection of AI-generated images. All of them rely on variants of CLIP, however only in [56] it is pursued generalization to OOD data, by nearest neighbor and linear probing in the CLIP feature space. To this end, a large dataset of fake and real images is used to train the classifier. In this work we make a further step in this direction and show that better performance can be achieved even using much less data.
@@ 3 Datasets and metrics
The main objective of this work is to study the generalization ability of CLIP-based and conventional methods, that is, assess how the detectors behave on data never seen before, of different provenance, and possibly impaired by various forms of post-processing. Therefore, we consider a very large test dataset and try to avoid any possible biases in the experimental validation.
@@
Tab. 1 lists all the generation models used in testing. They are grouped in three families: GAN-based (ProGAN, StyleGAN2, StyleGAN3, StyleGAN-T, GigaGAN), DM-based (Score-SDE, ADM, GLIDE, eDiff-I, Latent and Stable Diffusion, DiT, DeepFloyd-IF, Stable Diffusion XL) and commercial tools (DALL·E 2, DALL·E 3, Midjourney v5 and Adobe Firefly), which includes some recent text-driven methods available on-line, whose architecture is not always disclosed. These models cover the different generation modalities commonly adopted in this field: unconditional (u), conditional (c), and text-to-image (t), reported in second column of the table.
The last column reports the resolution of the images in the dataset. The central columns, instead, indicate which datasets of real images have been considered to carry out the detection experiments. These have been carefully selected to avoid all foreseeable biases. For example, when testing on synthetic images generated by ProGAN, we use as real counterparts images drawn from the LSUN dataset on which the ProGAN model was trained. Therefore, if LSUN images shared all the same defect (say, a red ”LSUN” label in a corner) the ProGAN images should spot the same defect, which could not be used to spoil detection performance. This hilarious example should not hide the importance of the problem, as biases are way too common in experiments and not always so easy to avoid. In other cases, for example when text-driven commercial tools with undisclosed training details are used, we relied on a dataset recently proposed in [4]. Pristine and synthetic images with similar semantic content are built by extracting textual descriptions from a real dataset (RAISE [16]) and then using them as prompt for generating synthetic ones. Overall, we selected $32,000$ real and fake images for testing, which will form the dataset used in all forthcoming experiments. To consider a more realistic scenario, we also consider datasets with post-processing operations that simulate images shared over social networks. In particular, we apply random cropping, resizing and JPEG compression at different quality levels. More details on the dataset can be found in the supplemental material, here we only show a few examples in Fig. 2.
@@
To measure performance, we consider three metrics. The area under the receiver-operating curve (AUC), and the average precision (AP), equal to the area under the precision-recall curve, are both threshold-independent. The Accuracy instead, given by the number of correct predictions divided by the number of tested images, depends on the decision threshold. We use always a fixed threshold equal to 0.5, which fits a realistic scenario with no prior information on the data under test to carry out proper calibration.
@@
Table 1: Image generators used in our experiments: GAN-based, DM-based, and commercial tools with models not always disclosed. For unconditional (u) and conditional (c) models, to prevent biases, the real images used in test come from the same datasets used to train the generator. Image resolution varies from 256$\times$256 to 2048$\times$2048.
@@ 4 CLIP for synthetic image detection
We propose a simple procedure to tell apart real images from synthetic images based on features extracted from the image encoder of CLIP ViT L/14. The design of the detector consists of the following four steps:
@@
    use the captions to feed a text-driven image generator, $G(\cdot)$, so as to obtain $N$ synthetic images $\{F_{1},\ldots,F_{N}\}$ with $F_{i}=G(t_{i})$. We have now $N$ real / fake pairs that share the same textual description;
@@
    use the $N$+$N$ feature vectors to design a linear SVM classifier.
If the real images have low-quality associated captions, or no caption at all, these can be easily generated by a dedicated tool such as BLIP [45]. Real and fake images are coupled to avoid possible semantic biases, which proves beneficial when the number of images is very small, e.g., $N=10$. We chose to extract features from the next-to-last layer rather than the last layer based on the results of preliminary experiments and, in the same way, we selected SVM among a set of candidate simple classifiers. All these preliminary experiments and their results are described in the supplemental material.
In the following, we analyze the performance of the proposed detector as a function of the main design choices and experimental conditions.

### v2

@@ Abstract
The aim of this work is to explore the potential of pre-trained vision-language models (VLMs) for universal detection of AI-generated images. We develop a lightweight detection strategy based on CLIP features and study its performance in a wide variety of challenging scenarios. We find that, contrary to previous beliefs, it is neither necessary nor convenient to use a large domain-specific dataset for training. On the contrary, by using only a handful of example images from a single generative model, a CLIP-based detector exhibits surprising generalization ability and high robustness across different architectures, including recent commercial tools such as Dalle-3, Midjourney v5, and Firefly. We match the state-of-the-art (SoTA) on in-distribution data and significantly improve upon it in terms of generalization to out-of-distribution data (+6% AUC) and robustness to impaired/laundered data (+13%). Our project is available at https://grip-unina.github.io/ClipBased-SyntheticImageDetection/
@@ 1 Introduction
Synthetic images have by now left research laboratories and are flooding the real world. The latest versions of popular image editing tools, such as Adobe Photoshop and Microsoft Paint, come with easy-to-use AI-powered generative tools that allow even novice users to edit and generate visual content at will. Thanks to diffusion-based generative models, it is not only the quality of the generated images that is greatly improved, but also the flexibility in using these tools. By issuing a few text commands you can easily obtain the desired image. While this represents a great opportunity for visual arts applications, it is also the paradise of disinformation professionals who can design their attacks with unprecedented power and flexibility [8, 5, 26]. And, of course, a nightmare for all those trying to combat the spread of fake news and orchestrated disinformation campaigns: journalists, fact-checkers, law enforcement, governments. Therefore, there is a high demand for automatic tools that help establish the authenticity of a media asset.
@@
Figure 1: Area Under ROC Curve (AUC %) on unseen synthetic generators ($x$-axis) and on post-processed data ($y$-axis). The first number measures the generalization ability of the detector, the second measures its robustness to possible impairments. Circle area is proportional to training set size. Performance is measured over 18 different synthetic models. Our CLIP-based detector largely outperforms all SoTA methods with very limited training data.
@@
It is well known that each synthesis model leaves its own peculiar traces in all generated images, subtle traces that give rise to so called artificial fingerprints, that can be exploited for forensic analyses [50, 79]. This phenomenon was first observed for methods based on generative adversarial networks (GAN) [81, 23] but holds, with obvious differences, for all generation approaches, including the more recent diffusion models (DM) [13]. One of the main challenges for current forensic detectors is the ability to generalize to new and unseen generative methods. Indeed, the in-distribution (ID) scenario, with perfectly aligned training and test sets, is rarely met in practice. Test images are often generated by new approaches, unseen architectures, or even known architectures re-trained under different conditions, all representing different flavors of the out-of-distribution (OOD) scenario. Additionally, most images are downloaded from social networks, where they undergo transformations such as compression and resizing (sometimes multiple times), processes that tend to wash out the tiny traces so precious for detection. In [73], it is shown that training diversity and intense augmentation are crucial for generalization. A ResNet-50 detector is trained on a single but highly diverse dataset of about 360k ProGAN [41] images. Results on OOD images generated by other GAN-based generators turned out to be surprisingly good. However, the authors themselves attribute this performance to the structural similarities between GAN-based generators. Indeed, results are not equally good on images generated by recent diffusion-based methods, which present somewhat different generation artifacts [14].
@@
Generalizing to OOD data is a major issue in deep learning and has been the object of intense research in recent years. In this context, the advent of large pre-trained vision-language models has brought about a number of new solutions and exciting results. These models have been shown to be excellent zero-shot and few-shot learners in many diverse applications, such as image classification [60, 80], detection [27, 54] and segmentation [76, 83]. Recently, there have been attempts to exploit the power of VLMs to detect synthetic images [67, 2, 57]. For example, [57] considers the same dataset, augmentation strategy and experimental protocol as in [73] but uses a pre-trained VLM, the Contrastive Language-Image Pre-Training (CLIP) [60], as a feature extractor rather than training a ResNet-50. Only the classifier is then learned on the task-specific dataset. Compared to [73] the performance improves significantly, especially on images generated by diffusion models, showing excellent generalization ability.
@@
In this work, we explore in depth the potential of CLIP for image forensics and conduct an extensive experimental analysis in challenging real-world scenarios involving a large number of generative models. We find that CLIP-based methods show impressive generalization ability and very good robustness, resulting in large performance improvements compared to the state of the art. To achieve such results, we avoid any intensive training on domain-specific data, which could introduce unwanted biases and undermine the descriptive power of CLIP features. Instead, we rely on a small number of paired real/fake images with the same textual description and use their CLIP features to model the decision space. Synthetic images come from a single generator, but results are equally good across all different data sources. The top performance is achieved with just 1,000 to 10,000 paired images. Moreover, only minor performance decays are observed when this number reduces to 100 or even 10 (see Fig. 1). In summary, the key contributions of this work are as follows:
@@
  • We carry out a large set of experiments on diverse synthetic generators and in very challenging conditions achieving the best performance on average. Our experiments make clear that the features extracted are partially orthogonal to the low-level features used by previous methods.
@@ 2 Related work
There is an extensive literature on synthetic image detection, primarily focusing on images created by generative adversarial networks (GANs) and, more recently, by diffusion models (DMs). Some methods look for visible errors, such as asymmetries in faces, incorrect perspectives, or unusual shadows [52, 29, 30]. However, with the rapid progress in image synthesis technology, these types of problems are solved quickly and appear less frequently in modern generation methods. Therefore, we will focus on methods that exploit “invisible” forensic traces, working either in the spatial or frequency domain. Subsequently, we will provide a brief overview of the most recent approaches that exploit multimodal features.
@@
Spatial domain methods. Despite their high visual quality, synthetic images carry with them distinctive traces of the generation process that enable detection and even pattern identification [50, 79]. Each generation model, in fact, inserts a sort of digital fingerprint into all the images it creates, which depends on its architectural and training details. This fingerprint can be easily estimated. Given a few hundred images generated by the target model, it is sufficient to average their noise residuals, extracted using simple denoisers [50] or more sophisticated methods based on deep learning [68, 48]. Detectors based on digital fingerprints can be regarded as “few-shot”, as they can deal with new models based on just a few example images. Other few-shot solutions have been proposed lately in this field [15, 51, 22, 37]. However, they all require information on the target models, even if limited to a few images, so they fit a strictly in-distribution scenario. In contrast, our work aims to generalize to new and previously unseen models, and therefore fits into an out-of-distribution scenario.
@@
Good generalization is a key requirement of synthetic image detectors. Towards this end, it is important to increase diversity in training, which can be pursued by suitable augmentation, by including a large number of different categories [73], or even by model ensembling [49]. Other studies suggest working on local patches [9] or combine global spatial information with local features [39]. In [34], it is shown that to preserve the subtle high-frequency forensic traces, one should avoid any downsampling in the first layers of the network. With the same aim, [70] proposes to work on noise residuals rather than the original images, extracting the gradients with a pre-trained CNN. In [12], a systematic study is carried out on transferable forensic features that allows easier generalization.
@@
The above investigations and proposals, however, only consider GAN-based generators. Some very recent works extend the analyses to latest generation approaches. [14] shows that detectors designed for GAN images have a hard time generalizing to DM images, especially in the presence of common post-processing steps such as compression and resizing. Furthermore, they are unable to select an adequate decision threshold without the help of some calibration data from the model under test. On the other hand, [25] shows that it is possible to achieve decent performance by continuously re-training the detector on images from new generators, as long as the latter share some architectural components with the known ones. In [74], inspired by previous work on GAN images [1], a detector specifically tailored to DM images is proposed. Images are projected in a latent space and reconstructed with known models to study their features. Although promising, the approach refers to specific architectures, hence an ID scenario, and has been tested on a limited set of generative diffusion models. In this work instead, we do not make any assumptions and include a large variety of both GAN and DM generators.
@@
Frequency domain methods. GAN-image artifacts are more easily spotted in the frequency domain and are clearly visible in the artificial fingerprint spectra [23, 32, 24]. In fact, they are caused by the up-sampling operations used in the generator, which give rise to regular spatial patterns and strong peaks in the Fourier domain. Interestingly, the awareness of such weakness has prompted the design of new architectures, like StyleGAN3 [44], which explicitly avoid aliasing and reduce the above-mentioned peaks. In addition to such obvious artifacts, the spectral content of GAN images and real images is also known to differ significantly in the medium and high frequencies. Likewise, clear differences in the radial and angular spectral distributions of the DM and real images were observed [13, 77]. Furthermore, it is worth observing that DM images may also present spectral peaks (see Fig. 5). Of course, frequency domain traces can be used to train simple detectors for ID images. Even more interestingly, some authors [81, 38] replicate and modify the synthesis process of known generators by introducing a series of small architectural changes to learn to handle even OOD test images. A major problem with spectral traces is their low resilience to laundering operations, both unintentional (e.g. image resizing) and intentional (e.g., counterforensic methods [21, 11, 16]).
@@
Methods based on multimodal features. The introduction of large language models has sparked intense work in the vision community attempting to leverage large models trained using both images and text [82, 80]. In image forensics, however, traditionally oriented towards the analysis of low-level features, there has been very little work in this direction. Only a few papers [67, 2, 57] have tried to leverage pre-trained visual models for AI-generated image detection. All these studies rely on variants of CLIP. However, only in [57] is generalization to OOD data pursued, via nearest neighbor and linear probing in CLIP feature space. To this end, a large dataset of fake and real images is used to train the classifier. In this work we take a further step in this direction and show that superior performance can be achieved even using much less data.
@@ 3 Datasets and metrics
The main objective of this work is to study the generalization capacity of conventional and CLIP-based methods, i.e. to evaluate how detectors behave on data never seen before, of different origins and possibly impaired by various forms of post-processing. Therefore, we consider a very large test dataset and take measures to mitigate any potential biases in the experimental validation process.
@@
Tab. 1 lists all generation models used during test. They are grouped into three families: GAN-based (ProGAN, StyleGAN2, StyleGAN3, StyleGAN-T, GigaGAN), DM-based (Score-SDE, ADM, GLIDE, eDiff-I, Latent and Stable Diffusion, DiT, DeepFloyd-IF, Stable Diffusion XL) and commercial tools (DALL·E 2, DALL·E 3, Midjourney v5 and Adobe Firefly), which include some recent text-based methods available online, whose architecture is not always disclosed. The second column of the table specifies the generation modalities employed by these models: unconditional (u), conditional (c), and text-to-image (t). The last column reports the resolution of the images in the dataset. The central columns, indicate which datasets of real images were considered to carry out the detection experiments. These have been carefully selected to avoid all foreseeable biases. For example, when testing synthetic images generated by ProGAN, we use as real counterparts images from the LSUN dataset on which the ProGAN model was trained.
The StyleGAN-T [64], eDiff-I [3], and GigaGAN [40] images were provided by the authors of the respective papers. Instead, we generated the images ourselves for Score-SDE [69], Stable Diff. [63], DiT [58], and Deepfloyd-IF [45] using the pre-trained models available online. Additional generated images were sourced from two public datasets [73, 14]. For the text-driven commercial tools, we used a recently proposed dataset [4]. Pristine and synthetic images with similar semantic content are built by extracting textual descriptions from a real dataset (RAISE [17]) and using them as prompts for generating synthetic images. A few examples of such images are shown in Fig. 2.
Overall, we have a dataset with $32,000$ real and fake images for all upcoming tests. To consider a more realistic scenario we simulate images shared on social networks and subject to post-processing operations: random cropping that can vary in a range from $\frac{5}{8}$ to the full size of the image, resizing to $200\times 200$ pixels, and JPEG compression with a random quality factor between $65$ and $100$.
@@
To measure performance, we consider three metrics. The area under the receiver operating curve (AUC) and average precision (AP) are both threshold-independent. The Accuracy instead, given by the number of correct predictions divided by the number of tested images, depends on the decision threshold. We always use a fixed threshold of 0.5, to simulate a realistic scenario in which no prior information on the data under test is available to carry out calibration.
@@
Table 1: Image generators used in our experiments: GAN-based, DM-based, and commercial tools. For unconditional (u) and conditional (c) models, the real images used in test come from the same datasets used to train the generator, while for text-to-image (t) models synthetic images have been generated using the prompt extracted from the real counterpart.
@@ 4 CLIP for synthetic image detection
We propose a simple procedure to distinguish real images from synthetic images based on features extracted from the image encoder of CLIP ViT L/14. The design of the detector consists of the following four steps:
@@
    use the captions to feed a text-driven image generator, $G(\cdot)$, so as to obtain $N$ synthetic images $\{F_{1},\ldots,F_{N}\}$ with $F_{i}=G(t_{i})$. Now we have $N$ real / fake pairs that share the same textual description;
@@
    use these $N$+$N$ vectors to design a linear SVM classifier.
If the real images have low-quality associated captions, or no caption at all, these can be generated by a dedicated tool such as BLIP [46]. Real and fake images are coupled to avoid possible semantic biases, which proves beneficial when the number of images is very small, e.g., $N=10$. We chose to extract features from the next-to-last layer rather than the last layer based on the results of preliminary experiments and, in the same way, we selected SVM among a set of candidate simple classifiers. All these preliminary experiments and their results are described in the supplementary material. In the following, we analyze the performance of the proposed detector as a function of the main design choices and experimental conditions.